Calculate the molarity of a solution prepared by dissolving 78.6g of KF in enough water to produce 225ml of solution. (at mass of F= 19g/mol, K= 39.10g/mol
1
Expert's answer
2012-12-24T10:44:00-0500
molarity = amount / liter

amount n = m/Mw
m-mass
Mw - molecular mass

in the question m = 78.6 g
Mw = 58.1 g/mol

n = 78.6/58.1 = 1.35 mol, but it is in 225 ml, and in 1000 ml or 1 L it would be X

X = 1.35*1000/225 = 6 M